What is the expected after-tax cash flow from selling a piece of equipment if Litchfield Design purchases the equipment today for 50,000 dollars, the tax rate is 40 percent, the equipment is sold in 3 years for 8,000 dollars, and MACRS depreciation is used where the depreciation rates in years 1, 2, 3, 4, and 5 are 30 percent, 27 percent, 25 percent, 11 percent, and 7 percent, respectively?
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=50,000*(1-0.3-0.27-0.25)

=9000

Hence loss on sale=9000-8000=$1000

Hence after-tax cash flow=Sale proceeds+(Tax rate*loss on sale)

=8000+(1000*40%)

=$8400